OpenStudy (faf3):

g(x)=2x-5 h(x)=6x write the expressions for (g+h)(x) and (g * h)(x) and evaluate (g-h)(2)

OpenStudy (anonymous):

(g+h)(x) = 2x-5+6x Simplify (g*h)(x) = (2x-5)(6x) Simplify (g-h)(x) = 2x-5-6x Simplify